OpenStudy (anonymous):

Integrate: (xˆ2+1)/2xˆ2

sam (.sam.):

\[\int\limits \frac{x^2+1}{2 x^2} \, dx\]

sam (.sam.):

Factor 1/2, \[\frac{1}{2} \int\limits \frac{x^2+1}{x^2} \, dx\]

sam (.sam.):

Expand out fractions \[\frac{1}{2} \int\limits \left(1+\frac{1}{x^2}\right) \, dx\]

sam (.sam.):

Split \[\frac{1}{2} \int\limits \frac{1}{x^2} \, dx+ \frac{1 }{2} \int\limits 1dx\]

sam (.sam.):

Integrate using power rule \[\int\limits x^n=\frac{x^{n+1}}{n+1}+c\]

OpenStudy (anonymous):

@.Sam. Thanks!!(: